OpenStudy (brydenwright):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 4>, v = <9, -9>

OpenStudy (sshayer):

\[u.v=\left| u \right|\left| v \right| \cos \theta \] \[where~ \theta~is~the~\angle~\between~u~and~v \]

OpenStudy (mathmale):

The quantity on the left side of this equation is the so-called DOT PRODUCT. I'd suggest you look this up and review it. In the same section you'll find a formula for the angle between two vectors. sshayer is correct. Can you take this info and come up with a solution now? Hint: solve sshayer's equation for cos theta first. find cost theta. then use the inverse cosine function to determine the angle, theta. Be careful: decide whether you want to work in degrees or in radians.

OpenStudy (sshayer):

u.v=(8)(9)+(4)(-9)=72-36=? \[\left| u \right|=\sqrt{8^2+4^2}=?\] similarly find |v| then proceed

OpenStudy (brydenwright):

so sqrt80 times sqrt162 cos x = 36?

OpenStudy (mathmale):

Please show all details of your calculations here. Where did that sqrt 80 come from? sqrt 162?